OpenStudy (anonymous):

Which of the following points does NOT lie on the graph of the following equation? 2x + -3y = -5

OpenStudy (mww):

Remember a point lies on the line if it satisfies the equation Substitute the y value and x value of the point in and see if the equation makes sense. If not, then the point does not lie on the line.

OpenStudy (anonymous):

A.(-4, -1) B. (2, 3) C. (4, 1) D. (8, 7) those are the choices, PLEASE HELP

OpenStudy (mww):

Do what I suggested. Take the point A(-4,1) as an example. In this case x = -4 and y =1 2x -3y = -5 is the equation. Substitute x = -4 and y =1 into it: Then consider LHS and RHS (left and right hand sides of equation) LHS = 2(-4) - 3(1) = -8 - 3 = -11 RHS = -5 but LHS is not equal to RHS. so this point does not lie on the line. When you find a point so that LHS = RHS, then it lies on the line.